OpenStudy (anonymous):

If .1% of m is equal to 10% of n, then m is what percent of 10n? How do I solve this algebraically?

OpenStudy (anonymous):

.1%m=10%n > .1/100m=10/100n > 10*1/100m= 10*10/100n continue

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

but the answer booklet says 1000%

OpenStudy (anonymous):

10*1/100m=10*10/100n > 1/10m=n > 10/10m=10n > m=10n 10*100%=1000%

OpenStudy (anonymous):

Ooh i get it thank you!